# A Class of Countably Covered Two-Dimensional Sofic Shifts

### v1

@@ 5 Proof of the main result
Next, we use Lemma 17 to add another copy of $X_{\mathrm{grid}}$, called the large grid layer, and couple its width to that of the small grid layer using the function $f+2$. If the small grid layer has a grid of some finite width $n\in\mathbb{N}$, now the large grid layer must have a grid of finite width $m\geq f(n)+2$. We choose $g(n)$ as the maximum value of $m$, the existence of which is guaranteed by Lemma 17.
Finally, we use Lemma 12 to add a counting layer that couples the large grid layer with the $A$-layer. It has the effect of constraining the number of nonzero columns on the $A$-layer to be at most $m-2$, but places no other restrictions on it.
@@
Let then $y\in Y$ be arbitrary, and let $z\in Z$ be such that $y=\pi(z)$. If the small grid layer of $z$ does not contain a finite-width grid, then $y$ contains at most one nonzero column, so $y\in G(g)^{\dagger}$. Suppose then that the small grid layer contains a grid of width $n$. Then Lemma 17 guarantees that the large grid layer contains a grid of width $m\leq g(n)$, and Lemma 12 states that there are at most $m-2$ nonzero columns in $y$. The width restriction layer ensures that the distance between any two such columns cannot be less than $n$. Hence $y\in G(g)^{\dagger}$.
@@
We now begin the second phase of the construction. This involves adding new layers to the SFT cover $Z$ to produce a new countable SFT $Z^{\prime}$ with $\pi(Z^{\prime})=X^{\dagger}$. We first add a new copy of $X_{\mathrm{grid}}$ called the simulation grid, on which we simulate a sequence of tile sets $(T_{n})_{n\in\mathbb{N}}$ as in Example 15; they will have access to the $A$-layer and are responsible for checking its correctness. Here, $n$ is the width of the small grid layer. The size of the simulation grid is at least $t(n)\cdot m$ for a computable function $t:\mathbb{N}\to\mathbb{N}$ growing fast enough that the counter machines always have enough time to correctly simulate tiles of $T_{n}$ (recall that $m$ is the size of the large grid); we enforce this using Lemma 17. We require that the small grid, large grid and simulation grid all be aligned, so that the counter machines running on the simulation grid cells can access $n$ and the contents of the counting layer.
@@
We define the simulated tile sets $T_{n}$ for $n\in\mathbb{N}$ as SFTs, which can then be recoded into sets of Wang tiles. The set $T_{n}$ contains $k_{n}=\binom{f(n)}{2}$ Wang tile layers, one for each two-element subset $\{a,b\}\subset\{0,1,\ldots,f(n)-1\}$; we identify these numbers with the nonzero columns of the $A$-layer. Exactly one of these layers is active at a time, and the others are inactive. For a pair of adjacent horizontal rows over $T_{n}$, the active layer must follow the adjacency rules of its layer, while the remaining $k_{n}-1$ layers must be identical across the two rows. The role of the active layer is passed on cyclically, so that on a height-$k_{n}$ horizontal strip of tiles, each layer is active on ecaxtly one row. Hence, each layer is “stretched” by a factor of $k_{n}$.
@@
Each of the $k_{n}$ layers consists of several sublayers. The first sublayer, called the computation sublayer, contains a simulation of a deterministic bidirectional counter machine $C_{n}$. Its purpose is to repeatedly probe the $A$-layer of $Z$ and determine whether it is in $X^{\dagger}$. The second sublayer, called the anchor sublayer, is a vertically constant layer over the alphabet $\{0,1,2\}$, in which the horizontal patterns $10$, $20$ and $21$ are forbidden. The positions of the $1$-symbols on this layer are controlled by the counter machine that performs the simulation of $T_{n}$. The idea is that the $1$-symbols on the anchor sublayer of the layer corresponding to the pair $\{a,b\}$ span the region delimited by the $a$th and $b$th nonzero columns of the $A$-layer, so that the counter machine of the computation sublayer can be initialized near the $a$th column and check the configuration between it and the $b$th column.
In addition to these sublayers, the $k_{n}$ layers share a common sublayer over the alphabet $B$ of Example 15, which is used to probe the $A$-layer. The counter machine of each layer has two special counters $c_{1}$ and $c_{2}$. The $1$-symbols of the common $B$-layer must always coincide with the portion of the computation cone that is between the ends of these counters on the active layer. When the zig-zag head of the active layer steps on the $A\times\{0,1\}$-symbol of the $B$-layer, it can store it as part of its state.
The machine $C_{n}$ behaves as follows. First, it determines the number $m$ of $1$-symbols on the anchor sublayer. Then, using the $B$-layer via its special counters $c_{1}$ and $c_{2}$, it probes for the width $h(n)$ of the simulation grid and then the full contents of the $A$-layer on the union of all the grid cells that simulate $1$-symbols on the anchor layer. In this way it receives a word $w\in A^{m\cdot h(n)}$. It checks whether $|w|_{\neq 0}\leq f(d)$, where $d$ is the minimum distance between two nonzero symbols in $w$, rejecting (and producing a tiling error) if this is not the case. Next, $C_{n}$ checks whether ${}^{\infty}0w0^{\infty}\in X$, which is possible since $X$ satisfies the countable cover conditions. If this holds, then $C_{n}$ halts and accepts. If ${}^{\infty}0w0^{\infty}\notin X$, then by compactness there exists a number $p\geq 0$ such that $0^{p}w0^{p}\notin\mathcal{L}(X)$, which we can also compute since $X$ is effectively closed. After this, the machine probes the $p$ symbols to the left and right of $w$ on the $A$-layer. If the result is $0^{p}w0^{p}$, then $C_{n}$ rejects; otherwise, it accepts.
@@
We now describe how the counter machines of the simulation grid determine the positions of the $1$-symbols of the anchor layer. The simulation grid is aligned with the small and large grids, and the counter machine in each cell of the simulation grid has access to the contents of them and the counting layer. In particular, it can read the size $n$ of the small grid, as well as the leftmost position of a dashed vertical line on the synchronization layer in its grid cell. Denote by $e$ the distance of this line from the west border of the grid cell. The machine checks that $e<f(n)$, rejecting if not. Then, on each layer $\{a,b\}$, on the anchor sublayer it places a 00-symbol if $e<a$, a $1$-symbol if $a\leq e\leq b$, and a $2$-symbol if $b<e$. In this way we guarantee that every configuration contains at most $f(n)$ nonzero columns, and that each pair of these columns has a corresponding simulated layer on which the space between them is highlighted with $1$-symbols.
@@
We add two more sublayers to each of the $k_{n}$ layers of $T_{n}$, the purpose of which is to periodically restart the computation of the machine $C_{n}$ on the computation layer. The third sublayer is a grid, which is coupled with the run of $1$-symbols on the second sublayer similarly to the width restriction layer: if $201^{m}$ 2 occurs in the second sublayer, the third sublayer must have a grid of side length at most $m$. The fourth sublayer is a grid that is coupled to the third sublayer using Lemma 17 with a computable function $h^{\prime}:\mathbb{N}\to\mathbb{N}$ such that the simulated counter machine $C_{n}$ will always take at most $h^{\prime}(m)$ rows of tiles to finish its computation if the number of $1$-symbols on the anchor layer is $m$; again, such a function clearly exists. A vertical line of the latter grid erases whatever computation was taking place on the computation sublayer, and initializes a new computation on each occurrence of $01$ on the anchor layer (of which there can be at most one).
@@
Finally, there may be symbols $a\in A_{0}$ such that ${}^{\infty}0a0^{\infty}\notin X$. For each such symbol, there necessarily exists $p\geq 0$ such that $0^{p}a0^{p}\notin\mathcal{L}(X)$, and we forbid this horizontal pattern from the $A$-layer of $Z^{\prime}$. This concludes the construction of $Z^{\prime}$.
@@
We first claim that $Z^{\prime}$ is countable. As $Z$ was already proved countable, it suffices to fix all the layers of $Z$ (the $A$-layer, the small and large grids, and their width restriction and counting layers) and show that the remaining layers have countably many choices. If the small grid layer contains no finite-size grid, then neither does the simulation grid layer, by the last claim on Lemma 17. Hence, it contains at most one computation cone of a counter machine, which has a countable number of possible computation paths, plus a countable number of choices for other degenerate signals. Suppose then that the small grid layer contains a grid of size $n$, so that the large grid layer contains a grid of some size $m\geq f(n)+2$. Then the simulation grid layer contains a grid of size at least $t(n)\cdot m$, on which a tiling of $T_{n}$ is simulated (since the counter machines in the grid cells can probe the value of $n$).
@@
On each of the $k_{n}$ layers of $T_{n}$, the anchor layer is completely determined by the previously described layers. If the anchor layer does not contain a finite run of $1$-symbols, then it does not contain a finite-width grid, and neither does the fourth sublayer by Lemma 17. Hence the simulation sublayer contains at most one position where a computation of $C_{n}$ is initialized, and since $C_{n}$ is deterministic, this constitutes a countable number of choices. On the other hand, if the anchor layer contains a pattern $201^{k}$ 2, then the third sublayer contains a grid of width at most $k$, and the fourth contains a grid of width at most $h^{\prime}(m)$. The latter grid periodically restarts the computation of $C_{n}$ on the same column. Hence the number of choices is again countable.
@@
We now prove $\pi(Z^{\prime})=X^{\dagger}$. Take any $x\in X^{\dagger}$ and use it as the $A$-layer. Since $x\in G(f)^{\dagger}$, we can safely fill all the layers of $Z$, in particular putting a grid of finite width $n$ on the small grid layer, equal to the size of the smallest gap on the $A$-layer, and one of some width $m$ on the large grid layer. We choose to fill the counting layer so that the dashed horizontal lines are directly above the horizontal grid lines on the left end of the configuration. On the simulation grid layer we can safely put a grid of finite width at least $t(n)\cdot m$ by Lemma 17. By our choice of the function $t$, the counter machines inside these grid cells have enough time to correctly simulate a tiling of $T_{n}$.
@@
Consider a layer of $T_{n}$ corresponding to a pair $\{a,b\}$. The anchor sublayer of this layer is determined by the previous layers. If it does not contain a finite run of $1$-symbols, then the remaining sublayers can safely be filled with degenerate, uniform configurations. If it does contain a run $201^{k}$ 2, we can put a grid of width $k$ on the third sublayer, and a grid of some width at least $h^{\prime}(k)$ on the fourth.
@@
This leaves the computation sublayer (and the common $B$-layer, which is determined by the previous layers and only constrains the computation sublayers). On this layer we are forced to initialize computations of the machine $C_{n}$ on the intersections of the horizontal grid lines of the fourth sublayer and the leftmost $1$-column on the anchor sublayer. The machine probes the small and large grid layers and the $A$-layer. Since the $A$-layer comes from $x\in X$, it cannot reject; the only way to do so would be to find too many nonzero columns or a forbidden word of the form $0^{p}w0^{p}$, neither of which can occur on the $A$-layer. Hence we have produced a valid tiling of $Z^{\prime}$.
@@
Conversely, take any $z\in Z^{\prime}$; we claim that its $A$-layer is in $X^{\dagger}$. If the small grid layer is degenerate, then so is the large grid layer, and the $A$-layer may contain at most one nonzero column. Since we explicitly forbade any ${}^{\infty}0a0^{\infty}\notin X$ from occurring on the $A$-layer, this case is handled. Suppose hence that the small grid layer contains a grid of size $n$. Then, as in the argument for the countability of $Z^{\prime}$, we must have finite-size grids on the large grid and simulation grid layers, the latter of which must simulate a tiling of $T_{n}$, which in turn consists of the $k_{n}$ layers plus the shared $B$-layer. The simulation layer also restricts the vertical dashed lines on the counting layer of $Z$ to be at height at most $f(n)$, and assigns to each nonzero column of the $A$-layer a unique successive number in $\{0,\ldots,f(n)\}$. Let these numbers span a sub-interval $\{a,a+1,\ldots,b\}$.
@@
We restrict our attention to the layer of the simulated $T_{n}$-tiling corresponding to the pair $\{a,b\}$. The run of $1$-symbols of its anchor sublayer is finite and spans the entire nonzero part of the $A$-layer. Thus, the third and fourth sublayers contain grids of finite size, and the computation sublayer contains an infinite number of restarted computations of $C_{n}$ that have enough time to complete by accepting. The machine $C_{n}$ reads the entire nonzero part of the $A$-layer as the word $w$. If we had ${}^{\infty}0w0^{\infty}\notin X$, then $C_{n}$ would have to find a nonzero column within a finite distance to the left or right of $w$, which is impossible, since $w$ spans all the nonzero columns. Hence we have ${}^{\infty}0w0^{\infty}\in X$, as required. ∎
6 Further results

### v2

@@ 5.1 Embedding $X^{\dagger}$ in a countably covered $G(g)^{\dagger}$
Next, we use Lemma 20 to add another copy of $X_{\mathrm{grid}}$, called the large grid layer, and couple its width to that of the small grid layer using the function $f+2$. If the small grid layer has a grid of some finite width $n\in\mathbb{N}$, now the large grid layer must have a grid of finite width $m\geq f(n)+2$. We choose $g(n)$ as the maximum value of $m$, the existence of which is guaranteed by Lemma 20; in other words, $g(n)=\max K$, where $K\subset[f(n)+2,\infty)$ is the finite set given by the Lemma.
Finally, we use Lemma 15 to add a counting layer that couples the large grid layer with the base alphabet layer. It has the effect of constraining the number of nonzero columns on the base alphabet layer to be at most m2m-2, but places no other restrictions on it.
5.2 Correctness of the embedding
@@
For the other inclusion, let $y\in Y$ be arbitrary, and let $z\in Z$ be such that $y=\pi(z)$. If the small grid layer of $z$ does not contain a finite-width grid, then $y$ contains at most one nonzero column, so $y\in G(g)^{\dagger}$. Suppose then that the small grid layer contains a grid of width $n$. Then Lemma 20 guarantees that the large grid layer contains a grid of width $m$ for some $m\in K$, which here implies $m\leq g(n)$ due to our choice of $g$. Lemma 15 states that there are at most m2m-2 nonzero columns in $y$. The width restriction layer ensures that the distance between any two such columns cannot be less than $n$. Hence $y\in G(g)^{\dagger}$.
5.3 The simulation layer
@@
We now begin the second phase of the construction. This involves adding new layers to the SFT cover $Z$ to produce a new countable SFT $Z^{\prime}$ with $\pi(Z^{\prime})=X^{\dagger}$. We first add a new copy of $X_{\mathrm{grid}}$ called the simulation grid, on which we simulate a sequence of tile sets $(T_{n})_{n\in\mathbb{N}}$ as in Example 18; they will have access to the base alphabet layer via the kind of interface defined in that example and are responsible for checking its correctness. We will constrain $n$ to be equal to the width of the small grid layer.
The size of the simulation grid is at least $t(n)\cdot m$ for a computable function $t:\mathbb{N}\to\mathbb{N}$ growing fast enough that the counter machines always have enough time to correctly simulate tiles of $T_{n}$ (recall that $m$ is the size of the large grid); we enforce this using Lemma 20. We require that the small grid, large grid and simulation grid all be aligned, so that the counter machines running on the simulation grid cells can access $n$ and the contents of the counting layer.
5.4 The simulated tile sets $T_{n}$
We define the simulated tile sets $T_{n}$ for $n\in\mathbb{N}$ as SFTs, which can then be recoded into sets of Wang tiles. The set $T_{n}$ contains $k_{n}=\binom{f(n)}{2}$ Wang tile layers, one layer $L_{n,a,b}$ for each two-element subset $\{a,b\}\subset\{1,\ldots,f(n)\}$ with $a<b$. We identify these numbers with the nonzero columns of the base alphabet layer as they are counted from left to right by the counting layer; the number of a nonzero column is the height of the shaded region on the counting layer at that position (see Figure 3). In particular, if there are less than $f(n)$ nonzero columns, the count may not begin from 1. The layer $L_{n,a,b}$ is responsible for probing the contents of the base alphabet layer between the $a$th and $b$th nonzero columns and verifying that this part does not contain forbidden patterns of $X$.
In addition to these layers, the set $T_{n}$ contains one probe interface layer over the alphabet $B$ of Example 18, which is used to probe the base alphabet layer. Each layer $L_{n,a,b}$ will have access to this common layer, and through it the contents of the base alphabet layer and the small grid. For this purpose, we add a probe implementation layer to $Z^{\prime}$ and couple it to the base alphabet and simulation layers as in Example 18.
@@ 5.4 The simulated tile sets $T_{n}$
The probing interface of Example 18 allows only one symbol per row to be probed by a simulated tiling. Hence, we cannot allow the $k_{n}$ layers of $T_{n}$ to access the base alphabet layer independently of each other. For this reason, exactly one of the layers $L_{n,a,b}$ is active on a given horizontal row, and the others are inactive. For a pair of adjacent horizontal rows over $T_{n}$, the layer that is active on the bottom row must follow the adjacency rules of its layer between these two rows, while the remaining $k_{n}-1$ layers must be identical across the two rows. The role of the active layer is passed on cyclically, so that on a height-$k_{n}$ horizontal strip of tiles, each layer is active on exactly one row. Hence, each layer is “stretched” vertically by a factor of $k_{n}$.
5.5 The layers $L_{n,a,b}$ of the simulated tilings
@@
Each of the $k_{n}$ layers $L_{n,a,b}$ of $T_{n}$ consists of several sublayers. The first sublayer, called the computation sublayer, contains a simulation of a deterministic bidirectional counter machine $C_{n}$, which is periodically restarted. Its purpose is to repeatedly probe the base alphabet layer of $Z$ between the $a$th and $b$th nonzero column and determine whether it is in $X^{\dagger}$.
The second sublayer, called the anchor sublayer, is a vertically constant layer over the alphabet $\{0,1,2\}$, in which the horizontal patterns $10$, 2020 and $21$ are forbidden. The machine $C_{n}$ has access to the anchor sublayer. We will force the $1$-symbols on the anchor sublayer of $L_{n,a,b}$ to span exactly the region delimited by the $a$th and $b$th nonzero columns of the base alphabet layer, so that the counter machine $C_{n}$ of the computation sublayer can be initialized near the $a$th column and check the configuration between it and the $b$th column.
@@
The positions of the $1$-symbols of the anchor layer are determined by the counter machines on the simulation grid as follows. The simulation grid is aligned with the small and large grids, and the counter machine in each cell of the simulation grid has access to the contents of them and the counting layer. In particular, it can read the width $n$ of the small grid, as well as the position of the westmost dashed vertical line on the synchronization layer in its grid cell. Denote by $e\geq 1$ the distance of the dashed line from the west border of the grid cell; recall that this distance is incremented between two adjacent small grid cells if and only if the western one contains a nonzero column. The machine checks that $e\leq f(n)$, rejecting if not. Then, on the anchor sublayer of each layer $L_{n,a,b}$ it places a 0-symbol if $e<a$, a $1$-symbol if $a\leq e\leq b$, and a $2$-symbol if $b<e$. In this way we guarantee that every configuration contains at most $f(n)$ nonzero columns, and that each pair of these columns has a corresponding simulated layer on which the space between them is highlighted with $1$-symbols.
5.6 The counter machine $C_{n}$ of $L_{n,a,b}$
Recall that the layer $L_{n,a,b}$, and hence both its anchor sublayer and the machine $C_{n}$, are part of the tiling over $T_{n}$ simulated on the width-$h(n)$ cells of the simulation grid. We now explain how the machine $C_{n}$ of $L_{n,a,b}$ accesses the contents of the base alphabet layer. It has two special counters $c_{1}$ and $c_{2}$. The (possibly empty) segment of $1$-symbols of the common probe interface layer of $T_{n}$ must always coincide with the portion of the computation cone of $C_{n}$ that is between the ends of these counters on the active layer. When the zig-zag head of the active layer steps on the $A\times\{0,1\}$-symbol of the probe interface layer, it can store it as part of its state.
The machine $C_{n}$ behaves as follows. First, it determines the number $m$ of $1$-symbols on the anchor sublayer of $L_{n,a,b}$. Then, using the probe interface layer via its special counters $c_{1}$ and $c_{2}$, it probes for the width $h(n)$ of the simulation grid [^2] Note that $h(n)$ is analogous to the width $m$ in Example 18, where we explain how the simulated tile set can obtain it using the probe interface layer. and then the full contents of the base alphabet layer on each simulation grid cell that simulates a $1$-symbol on the anchor layer. In this way it receives a word $w\in A^{m\cdot h(n)}$. It checks whether $|w|_{\neq 0}\leq f(d)$, where $d$ is the minimum distance between two nonzero symbols in $w$, rejecting (and producing a tiling error) if this is not the case.
Next, $C_{n}$ checks whether $^{\infty}0w0^{\infty}\in X$, which is possible since $X$ satisfies the countable cover conditions. If this holds, then $C_{n}$ halts and accepts. If $^{\infty}0w0^{\infty}\notin X$, then by compactness there exists a number $p\geq 0$ such that $0^{p}w0^{p}\notin\mathcal{L}(X)$, which we can also compute since $X$ is effectively closed. After this, the machine probes the $p$ symbols to the left and right of $w$ on the base alphabet layer. If the result is $0^{p}w0^{p}$, then $C_{n}$ rejects and produces a tiling error; otherwise, it accepts.
5.7 Restarting the machine $C_{n}$: the restart grids
We add two more sublayers to each of the $k_{n}$ layers of $T_{n}$, called the small and large restart grids, the purpose of which is to periodically restart the computation of the machine $C_{n}$ on the computation layer. Namely, if we allowed the machine to run indefinitely, then by shifting vertically we would obtain limit configurations where no computation is ever initialized.
@@ 5.7 Restarting the machine $C_{n}$: the restart grids
As the names suggest, both sublayers are copied of the grid shift $X_{\mathrm{grid}}$. The small restart grid is coupled with the run of $1$-symbols on the anchor sublayer similarly to the width restriction layer: if $01^{m}2$ occurs on the anchor sublayer, the small restart grid must have a grid of width between $1$ and $m$. The large restart grid is coupled to the small restart grid using Lemma 20 with a computable function $h^{\prime}:\mathbb{N}\to\mathbb{N}$ such that the simulated counter machine $C_{n}$ will always take at most $h^{\prime}(m)$ rows of tiles to finish its computation if the number of $1$-symbols on the anchor layer is $m$; again, such a function clearly exists. A vertical line of the large restart grid erases whatever computation was taking place on the computation sublayer, and initializes a new computation on each occurrence of the word $01$ on the anchor layer (of which there can be at most one).
@@
Finally, there may be symbols $a\in A_{0}$ such that $^{\infty}0a0^{\infty}\notin X$. These cannot be completely handled like the words $w$ such that $^{\infty}0w0^{\infty}\notin X$ at the end of Section 5.6, since they may occur as rows of the base alphabet layer in configurations without a finite-width small grid (and hence without simulations of any $T_{n}$ and $C_{n}$). For each such symbol, there necessarily exists $p\geq 0$ such that $0^{p}a0^{p}\notin\mathcal{L}(X)$, and we explicitly forbid this horizontal pattern from the base alphabet layer of $Z^{\prime}$. This concludes the construction of $Z^{\prime}$.
5.8 The SFT $Z^{\prime}$ is countable
We claim that $Z^{\prime}$ is countable. As $Z$ was already proved countable, it suffices to fix all the layers of $Z$ (the base alphabet layer, the small and large grids, and their width restriction and counting layers) and show that the remaining layers have countably many choices.
@@ 5.8 The SFT $Z^{\prime}$ is countable
If the small grid layer contains no finite-width grid, then neither does the simulation grid layer, by the last claim on Lemma 20. Hence, it contains at most one computation cone of a counter machine, which has a countable number of possible computation paths, plus a countable number of choices for other degenerate signals. Suppose then that the small grid layer contains a grid of width $n$, so that the large grid layer contains a grid of some width $m\geq f(n)+2$. Then the simulation grid layer contains a grid of width at least $t(n)\cdot m$, on which a tiling of $T_{n}$ is simulated (since the counter machines in the grid cells can access the value of $n$).
@@
On each of the $k_{n}$ layers $L_{n,a,b}$ of $T_{n}$, the anchor layer is completely determined by the previously described layers. If the anchor layer does not contain a finite run of $1$-symbols, then the small restart grid does not contain a finite-width grid, and neither does the large restart grid by Lemma 20. Hence the simulation sublayer contains at most one position where a computation of $C_{n}$ is initialized, and since $C_{n}$ is deterministic, this (together with degenerate computation cones) constitutes a countable number of choices. On the other hand, if the anchor layer contains a pattern $01^{k}2$, then the small restart grid contains a grid of width at most $k$, and the large restart grid contains a grid of width at most $h^{\prime}(m)$. The latter grid periodically restarts the computation of $C_{n}$ on the same column. Hence the number of choices is again countable.
5.9 The SFT $Z^{\prime}$ projects onto $X^{\dagger}$
@@
We now prove $\pi(Z^{\prime})=X^{\dagger}$. Take any $x\in X^{\dagger}$ and use it as the base alphabet layer. Since $x\in G(f)^{\dagger}$, we can safely fill all the layers of $Z$, in particular putting a grid of finite width $n$ on the small grid layer, equal to the size of the smallest gap on the base alphabet layer, and one of some width $m$ on the large grid layer. We choose to fill the counting layer so that the dashed horizontal lines are directly above the horizontal grid lines on the left end of the configuration. On the simulation grid layer we can safely put a grid of finite width at least $t(n)\cdot m$ by Lemma 20. By our choice of the function $t$, the counter machines inside these grid cells have enough time to correctly simulate a tiling of $T_{n}$.
@@
Consider a layer $L_{n,a,b}$ of $T_{n}$. The anchor sublayer of $L_{n,a,b}$ is determined by the previous layers. If it does not contain a finite run of $1$-symbols, then the remaining sublayers can safely be filled with degenerate, uniform configurations. If it does contain a run $01^{k}2$, we can choose a grid of width $k$ as the small restart grid, and a grid of some width at least $h^{\prime}(k)$ as the large restart grid.
@@
This leaves the computation sublayer (and the common probe interface layer, which is determined by the previous layers and only constrains the computation sublayers). On this sublayer we are forced to initialize computations of the machine $C_{n}$ on the intersections of the horizontal grid lines of the large restart grid and the leftmost $1$-column on the anchor sublayer. Since the base alphabet layer comes from $x\in X$, the machine $C_{n}$ cannot reject; the only way to do so would be to find too many nonzero columns or a forbidden word of the form $0^{p}w0^{p}$, neither of which can occur on the base alphabet layer. Hence we have produced a valid tiling of $Z^{\prime}$.
@@
Conversely, take any $z\in Z^{\prime}$; we claim that its base alphabet layer is in $X^{\dagger}$. If the small grid layer is degenerate, then so is the large grid layer, and the base alphabet layer may contain at most one nonzero column. Since we explicitly forbade any $^{\infty}0a0^{\infty}\notin X$ from occurring on the base alphabet layer, this case is handled.
Suppose hence that the small grid layer contains a grid of some width $n$. Then, as in the argument for the countability of $Z^{\prime}$, we must have finite-size grids on the large grid and simulation grid layers, the latter of which must simulate a tiling of $T_{n}$, which in turn consists of the $k_{n}$ layers $L_{n,a,b}$ plus the shared probe interface layer. The simulation layer also restricts the vertical dashed lines on the counting layer of $Z$ to be at height at most $f(n)$, and assigns to each nonzero column of the base alphabet layer a unique successive number in $\{0,\ldots,f(n)\}$. Let these numbers span a sub-interval $\{a,a+1,\ldots,b\}$.
@@
We restrict our attention to the layer $L_{n,a,b}$ of the simulated $T_{n}$-tiling. The run of $1$-symbols of its anchor sublayer is finite and spans the entire nonzero part of the base alphabet layer. Thus, the restart grids contain grids of finite size, and the computation sublayer contains an infinite number of restarted computations of $C_{n}$ that have enough time to complete by accepting. The machine $C_{n}$ reads the entire nonzero part of the base alphabet layer as the word $w$. If we had $^{\infty}0w0^{\infty}\notin X$, then $C_{n}$ would have to find a nonzero column within a finite distance to the left or right of $w$, which is impossible, since $w$ spans all the nonzero columns. Hence we have $^{\infty}0w0^{\infty}\in X$, as required. This finishes the proof of Theorem 6.
5.10 A modification